OpenStudy (anonymous):

if the equation x^2-x-n=0 has integer roots, the number of values of between 1 and 100 is

OpenStudy (anonymous):

Are you looking for all possible values for n?

OpenStudy (anonymous):

If so, n has to be the result of two numbers multiplied together that when added give you -1. This means the numbers have to be consecutive: n=1*2 n=2*3 n=3*4 n=4*5 n=5*6 n=6*7 n=7*8 n=8*9 n=9*10

OpenStudy (anonymous):

Those are all of them because if you went higher then n would be over 100 and if you went lower n would be less than 1

OpenStudy (anonymous):

u are right t bates